order: cipro 250 mg IVP q 12 h for urinary tract infection

on hand: cipro 200 mg in 20ml

how many milliliters will you add to the IV solution?
Answer #1

ANSWER

* Calculation of volume(ml) of drug required.

   Formula = Desired dose/Dose in hand * Quantity.

Here

    Desired dose(D) = 250 mg.

    Dose in hand (H) = 200 mg

        Quantity (Q) = 20 ml.

   As per above formula

    Volume(ml) of drug required = D/H * Q

                                                   = 250/200 * 20

                                                   = 1.25 * 20

                                                   = 25 ml.